OpenStudy (anonymous):

can some one help me with expansion and factorization of algebraic expression

OpenStudy (anonymous):

3(2a+b) how do i exspand this?

Directrix (directrix):

The distributive property of multiplication "distributes" the multiplication of 2 by (2a +b) over the "+". 3(2a+b) = 3 (2a) + 3(b) = 6a + 3b

OpenStudy (anonymous):

thank u now this one i have to exspand can u break it down in more steps for me plsss

OpenStudy (anonymous):

4(3X+4y)

Directrix (directrix):

4(3X+4y) = 4(3x) + 4(4y) = 12x + 16y

OpenStudy (anonymous):

now exsplain the steps u did to get the answer bc thats what im having trouble on

hero (hero):

It's called the distributive property. The distributive property states that for any integer a, b, and c a(b + c) = ab + ac For example 4(x + y) = 4x + 4y 3(2x + 7y) = 6x + 21y

OpenStudy (anonymous):

so your manly taking away the brackets?

OpenStudy (anonymous):

yes... try to convince yourself... is 4*(3+2) = 4*3 + 4*2 ??? 4*(3+2) = 4*(5) = 20 4*3 + 4*2 = 12 + 8 = 20

hero (hero):

You have to understand the purpose of the brackets...Think order of operations together with the distributive rule

OpenStudy (anonymous):

now whats the purpose of the brackets

OpenStudy (anonymous):

and whats the distributive rule ?

OpenStudy (anonymous):

it means that a(b + c) and a*b + a*c have the same VALUE, no matter what a, b, and c are.

OpenStudy (anonymous):

notice one is written with parenthesis and the other is not. they both have equal values.

OpenStudy (anonymous):

now is that for the distruibutive thing or for the purpose of brackets

OpenStudy (anonymous):

this EQUATION is called the distributive property of multiplication over addition: a(b+c) = ab + ac

OpenStudy (anonymous):

let's try some numbers to illustrate this... pick any 3 numbers...

OpenStudy (anonymous):

something you can calculate quickly would be preferable..

OpenStudy (anonymous):

486

OpenStudy (anonymous):

4, 8, and 6?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok when you see this expression 8(4+6) you would calculate what's in the parenthesis first to get 8(10) which is...

OpenStudy (anonymous):

4+6=10

OpenStudy (anonymous):

ok so what's 8(10) =

OpenStudy (anonymous):

that means 8 times 10

OpenStudy (anonymous):

80 thats if u mean whats 8*10

OpenStudy (anonymous):

yes, 8(10) means 8*10

OpenStudy (anonymous):

so, 8(4+6) = 80 now distributive property says we should get the same answer if we did this: 8*4 + 8*6

OpenStudy (anonymous):

8*4 = 8*6 =

OpenStudy (anonymous):

32 and 48

OpenStudy (anonymous):

ok, so what's 32 + 48 =

OpenStudy (anonymous):

80

OpenStudy (anonymous):

so, do you agree that 8(4 + 6) = 8*4 + 8*6 ? this is the distributive property... you need to know this property very well in algebra...

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

now can u explain statistics

OpenStudy (anonymous):

i'm glad you stuck with it asking questions. good work...

OpenStudy (anonymous):

i'll try... btw, did you know that 98% of all statistics are made up?

OpenStudy (anonymous):

i'm not that good at stats.

OpenStudy (anonymous):

r u serious i never knew that this is my first time working on statistics i tried youtbe vidoes this stff is hard to learn and im only in 8 grade and they want me to learn this stuff lmfao

OpenStudy (anonymous):

I'm part of that 98%... it was a joke, get it?

OpenStudy (anonymous):

no lets just study the statistics stuff lmfao

OpenStudy (anonymous):

anyway, if you have a different question please post it so other answerers can attempt at it... someone better at stats. ok?

OpenStudy (anonymous):

good luck...